# The foot, the fan, and the cuprate phase diagram: Fermi-volume-changing quantum phase transitions

### v7

@@ 5 Confinement crossover
Using fermion-boson duality, the $\pi$-flux spin liquid was argued by Wang et al. [142] to have dual description as the $\mathbb{CP}^{1}$ theory of bosonic spinons [143]. This bosonic spinon description is obtained by the complementary representation of fractionalization by a rotating reference frame in spin space [74, 75, 76, 77]
@@
It is worth mentioning here an important low $T$ difference between FL* and the holon metal which could be significant for transport experiments. As the holons are spinless, the area enclosed by a holon pocket is twice that of the hole pocket in FL* [146], and the holon pocket area is the same as that of the hole pocket in the ordered antiferromagnet. Very recent observations of the Yamaji effect [147] in the single layer cuprate HgBa2CuO4+δ at doping $p=0.1$ support the FL* theory of the pseudogap: there are 4 hole pockets of spin-1/2 fermions in FL*, and so each hole pocket occupies a fraction $p/8=0.0125$ of the Brillouin zone, to be compared to the observed value $\approx 1.3$% [147].
@@
    For a suitable Higgs potential [136], the condensation of $B$ leads to $d$-wave superconductivity, with four nodal fermionic quasiparticles. These quasiparticles have anisotropic velocities, similar to those in the BCS $d$-wave superconductor [140].

### v8

@@ 5 Confinement crossover
Using fermion-boson duality, the $\pi$-flux spin liquid was argued by Wang et al. [142] to have dual description as the $\mathbb{CP}^{1}$ theory of bosonic spinons [143, 144]. This bosonic spinon description is obtained by the complementary representation of fractionalization by a rotating reference frame in spin space [74, 75, 76, 77]
@@
It is worth mentioning here an important low $T$ difference between FL* and the holon metal which could be significant for transport experiments. As the holons are spinless, the area enclosed by a holon pocket is twice that of the hole pocket in FL* [147], and the holon pocket area is the same as that of the hole pocket in the ordered antiferromagnet (see Fig. 15).
Figure 15: From Kaul et al. [147], with added text on top. Fermi surfaces in the square lattice Brillouin zone obtained by doping an insulating antiferromagnet in the vicinity of a quantum phase transition from the Néel state to the valence bond solid (VBS). The Néel state has 2 pockets of spinful quasiparticles in the magnetic Brillouin zone, and so each hole pocket each occupies a fractional area $p/4$. (The holon metal has 4 pockets of spinless quasiparticles, and so the fractional area of each pocket remains $p/4$). The VBS order appears at long distances in the insulator above an underlying $\pi$-flux spin liquid [143, 144, 142], and the ultimate broken translational symmetry plays no role in Fermi surface structure in the computation in Kaul et al. [147]. Consequently the right panel applies also to a FL* metal: there are 4 pockets of spinful quasiparticles, and so the fractional area of each pocket is $p/8$, as stated by Kaul et al. [147], and close to the value observed by Chan et al. [148].
Very recent observations of the Yamaji effect [148] in the single layer cuprate HgBa2CuO4+δ at doping $p=0.1$ support the FL* theory of the pseudogap: there are 4 hole pockets of spin-1/2 fermions in FL*, and so each hole pocket occupies a fraction $p/8=0.0125$ of the Brillouin zone, to be compared to the observed value $\approx 1.3$% [148].
@@
    For a suitable Higgs potential [136], the condensation of $B$ leads to $d$-wave superconductivity, with four nodal fermionic quasiparticles. These quasiparticles have anisotropic velocities, similar to those in the BCS $d$-wave superconductor [149, 140].